Consider a step-index fiber for which n1 = 1.5, n2 = 1.30 and a = 25µm. What is the maximum value of numerical aperture.


1
Expert's answer
2021-05-09T15:11:57-0400

Given, "n_1=1.5,n_2=1.30"

The numerical aperture is-


"NA=\sqrt{n_1^2-n_2^2}"


"=\sqrt{(1.51)^2-(1.49)^2}=\sqrt{2.28-2.22}=\sqrt{0.06}=0.245"